# Lightweight and Generalizable Acoustic Scene Representations via Contrastive Fine-Tuning and Distillation

### v1

@@ 2 Contrastive Fine-Tuning
Model Architecture. We employ the pre-trained BEATs model [11] as our teacher backbone, which learns bidirectional encoder representations from audio transformers through iterative acoustic tokenization. Unlike prior ASC works [12, 9] that utilize PaSST as the teacher model designed for audio sampled at 32 kHz, BEATs operates on 16 kHz audio, making it more suitable for knowledge transfer to compact edge models that typically process audio at this sampling rate. The BEATs encoder produces 768-dimensional embeddings ($d=768$) that capture rich semantic audio representations.
@@
Cosine Classification Head. In our proposed method, we replace the standard linear classification head with a cosine similarity-based classifier to enhance embedding transferability. The cosine classifier computes logit for each class as $\text{logits}_{c}=\gamma\,\text{CSF}(\mathbf{x},\mathbf{w}_{c})$, where $\text{CSF}(\mathbf{p},\mathbf{q})=\mathbf{p}\cdot\mathbf{q}/\left(\|\mathbf{p}\|\|\mathbf{q}\|\right)$ is the cosine similarity function, $\mathbf{x}$ is the input embedding, $\mathbf{w}_{c}$ is the trainable weight vector for class $c$, and $\gamma$ is the cosine scale parameter. In practice, both $\mathbf{x}$ and $\mathbf{w}_{c}$ are $\ell_{2}$-normalized, simplifying the computation to scaled dot products. This formulation bounds the logits and reduces sensitivity to embedding magnitudes, promoting more generalizable feature representations [16]. We empirically set $\gamma=56$ based on the validation performance.
@@ 3 Contrastive Representation Distillation
Data Augmentation. In addition to the augmentations described in Section 2, we incorporate two further techniques for CRD: (i) device impulse response (DIR) augmentation, applied with probability 0.6 [21], and (ii) frequency shift augmentation, which randomly shifts the STFT’s maximum frequency range within $\pm$1000 Hz to simulate variations in frequency response. We omit mixup, as CRD requires well-defined positive and negative pairs.
@@
We set $\alpha=0.02$, $\beta=0.1$, knowledge distillation temperature $\tau_{\text{KD}}=2.0$, and the CRD temperature $\tau_{\text{CRD}}=0.07$. Following the hyperparameters set in CP-Mobile, we train five model variants (6K-126 K parameters) using AdamW optimizer with peak learning rates $\{0.04,0.04,0.03,0.02,0.01\}$, and cosine annealing scheduling over 75 epochs with 7 warmup epochs.
@@ 4 Results
Transferability Analysis. Table 3 provides insight into the transferability by decomposing the performance on seen and unseen acoustic classes. Compared to standard pipeline (FT + KD), our approach (C-FT + CRD) achieves comparable performance on seen classes, while demonstrates superior generalization to unseen categories. This highlights the effectiveness of ContrastASC for new category adaptation.

### v2

@@ 2 Contrastive Fine-Tuning
Model Architecture. We employ the pre-trained BEATs model [2] as our teacher backbone, which learns bidirectional encoder representations from audio transformers through iterative acoustic tokenization. Unlike prior ASC works [16, 14] that utilize PaSST as the teacher model designed for audio sampled at 32 kHz, BEATs operates on 16 kHz audio, making it more suitable for knowledge transfer to compact edge models that typically process audio at this sampling rate. The BEATs encoder produces 768-dimensional embeddings that capture rich semantic audio representations.
@@
Cosine Classification Head. In our proposed method, we replace the standard linear classification head with a cosine similarity-based classifier to enhance embedding transferability. The cosine classifier computes logit for each class as $\text{logits}_{c}=\gamma\,\text{COS}(\mathbf{x},\mathbf{w}_{c})$, where $\text{COS}(\mathbf{x},\mathbf{w}_{c})$ is the cosine similarity between the the input embedding $\mathbf{x}$ and the trainable weight vector $\mathbf{w}_{c}$ for class $c$, and $\gamma$ is the scale parameter. This formulation bounds the logits and reduces sensitivity to embedding magnitudes, promoting more generalizable feature representations [10]. We empirically set $\gamma=56$ based on the validation performance.
@@ 3 Contrastive Representation Distillation
Data Augmentation. In addition to the augmentations described in Section 2, we incorporate two further techniques for CRD: (i) device impulse response (DIR) augmentation [13, 21], applied with probability 0.6, and (ii) frequency shift augmentation, which randomly shifts the STFT’s maximum frequency range within $\pm$1000 Hz to simulate variations in frequency response. We omit mixup, as CRD requires well-defined positive and negative pairs.
@@
where $\mathcal{L}_{\text{KD}}$ is the standard knowledge distilation loss based on KL-divergence. We set $\alpha=0.02$, $\beta=0.1$, knowledge distillation temperature $\tau_{\text{KD}}=2.0$, and the CRD temperature $\tau_{\text{CRD}}=0.07$. Following the hyperparameters set in CP-Mobile, we train five model variants (6K-126K parameters) using AdamW optimizer with peak learning rates $\{0.04,0.04,0.03,0.02,0.01\}$, and cosine annealing scheduling over 75 epochs with 7 warmup epochs.
@@ 4 Results
Transferability Analysis. Table 3 provides insight into the transferability by decomposing the performance on seen and unseen classes. Compared to standard pipeline (FT+KD), our approach (C-FT+CRD) achieves comparable performance on seen classes, while demonstrates superior generalization to unseen categories (3/10 and 11/15 classes are unseen in TUT17 and ICME24, respectively). This highlights the effectiveness of ContrastASC for new category adaptation.